OpenStudy (anonymous):

The meat department in a grocery store needs 200 lb of hamburger that is 80% lean beef. How many pounds of 78% lean and 85% lean must be mixed?

OpenStudy (anonymous):

x+y=200 .78x + .85y = 200(.80) .78x+.85(200-x)=160 -.07x+170=160 -.07x=-10 x=142.86 78% beef y=57.14 85% beef